# Simple and accurate complete elliptic integrals for the full range of modulus

### v5

@@ Abstract
Keywords: computational physics, nonlinear pendulum, complete elliptic integrals, mathematical inequalities
@@ 1 Introduction
The complete elliptic integral of the first kind, denoted $K(k)$, is defined as
$K(k)=\int_{0}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{1-k^{2}\sin^{2}\theta}}$ (1)
and of the second kind,
$E(k)=\int_{0}^{\frac{\pi}{2}}\sqrt{1-k^{2}\sin^{2}\theta}\,d\theta$ (2)
@@ 2 Results
$E(k)\approx 1+\frac{1-k^{2}}{2n}\ln[\Big{(}\frac{4/\!\sqrt{e}}{\sqrt{1-k^{2}}}\Big{)}^{n}+b]$ (5)
@@
Figure 2: Percent error of $K(k)$ and $E(k)$ from various simple analytical forms.
@@
Figure 3: Error of $K(k)$ and $E(k)$ from various mathematical inequalities.
Fig. 3 (left) shows the performance of the upper and lower bounds for the complete elliptic integral $K(k)$. By resolving conjectures on inequalities between the $K(k)$ and simple logarithmic expressions, Qiu and coauthors derived sharp two-sided bounds for $K(k)$ [17]. These bounds are given in terms of elementary functions (notably logarithms in $k^{\prime}=\sqrt{1-k^{2}}$) and are asymptotically exact as $k\to 0$ and as $k\to 1$.
Fig. 3 (right) illustrates the sharp bounds for the complete elliptic integral $E(k)$. Yang and coauthors derived sharp lower and upper bounds for $E(k)$ in terms of elementary functions [20, 22]. The upper bound is also exact in the two limiting cases. In the inset, the approximation $E(k)$ in this work intersects the upper bound at $k\approx 0.964$. This crossover suggests an interesting property that may merit further investigation within the framework of inequality analysis.
@@ 3.1 Nonlinear Pendulum
Fig. 4 shows the phase space of an energy-conserving nonlinear pendulum with physical parameters $\sqrt{L/g}=1\,\text{second}$. The $\theta(t)$ are as follow.
$\omega_{m}<\omega_{c}:$ $\quad\theta(t)=\sum_{n\;\text{odd}}a_{n}\sin(n\Omega_{0}t+n\delta)$ (6)
$\omega_{m}=\omega_{c}:$ $\quad\theta(t)=2\arcsin\big{[}\tanh(\sqrt{\tfrac{g}{L}}t+\delta)\,\big{]}$ (7)
$\omega_{m}>\omega_{c}:$ $\quad\theta(t)=\Omega_{0}t+\delta+\sum_{n\geq 1}b_{n}\sin(n\Omega_{0}t+n\delta)$ (8)
@@
Specifically, the period $T$ for both classes of motion are, ($k=\sqrt{L/g}\,\omega_{m}/2$)
Swinging: $\quad T=\sqrt{\frac{L}{g}}4K(k)$ (9)
Spinning: $\quad T=\sqrt{\frac{L}{g}}\frac{2}{k}K(1/k)$ (10)
@@ 4 Theoretical Method
Complete asymptotic behaviors of $K(k)$ and $E(k)$ are rarely available in mathematical handbooks. Fortunately, they are documented in the NIST Digital Library of Mathematical Functions (DLMF) [29] and serve as the foundation for the proposed formulae in this work. The complete asymptotic behaviors are [29, Eq.19.12.1-2]
$K(k)$ $=\sum_{m=0}^{\infty}\frac{{(\frac{1}{2})}_{m}{(\frac{1}{2})}_{m}}{m!m!}{k^{\prime}}^{2m}\Big{[}\ln(\frac{1}{k^{\prime}})+d(m)\Big{]}$ (12)
$E(k)$ $=1+\frac{1}{2}\sum_{m=0}^{\infty}\frac{{(\frac{1}{2})}_{m}{(\frac{3}{2})}_{m}}{{(2)}_{m}m!}{k^{\prime}}^{2m+2}\Big{[}\ln(\frac{1}{k^{\prime}})+d(m)-\frac{1}{(2m+1)(2m+2)}\Big{]}$ (13)
where $k^{\prime}=\sqrt{1-k^{2}}$. The parenthesis with subscript, for example ${(\frac{1}{2})}_{m}$, is Pochhammer’s symbol. The auxiliary function $d(m)$ has a recursive relation $d(m+1)=d(m)-\frac{2}{(2m+1)(2m+2)}$, and $d(0)=\ln 4$. To find a balance between simplicity and accuracy, only the first term $m=0$ is kept. They become
$\text{Asymptotic Limit,}\qquad K(k)$ $\approx\ln\Big{[}\frac{4}{\sqrt{1-k^{2}}}\Big{]}$ (14)
$\text{Asymptotic Limit,}\qquad E(k)$ $\approx 1+\frac{1-k^{2}}{2}\ln\Big{[}\frac{4/\!\sqrt{e}}{\sqrt{1-k^{2}}}\Big{]}$ (15)
@@
The following form is proposed.
$K(k)\approx\frac{1}{n}\ln[\Big{(}\frac{4}{\sqrt{1-k^{2}}}\Big{)}^{n}+b]$ (16)
@@
The following form is proposed.
$E(k)\approx 1+\frac{1-k^{2}}{2n}\ln[\Big{(}\frac{4/\!\sqrt{e}}{\sqrt{1-k^{2}}}\Big{)}^{n}+b]$ (17)

### v6

@@ Abstract
Keywords: computational physics, nonlinear pendulum, complete elliptic integral, mathematical inequality
@@ 1 Introduction
The complete elliptic integral of the first kind, denoted $K(k)$, is defined as:
$K(k)=\int_{0}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{1-k^{2}\sin^{2}\theta}},$ (1)
and that of the second kind:
$E(k)=\int_{0}^{\frac{\pi}{2}}\sqrt{1-k^{2}\sin^{2}\theta}\,d\theta.$ (2)
@@ 2 Results
$E(k)\approx 1+\frac{1-k^{2}}{2n}\ln[\Big(\frac{4/\!\sqrt{e}}{\sqrt{1-k^{2}}}\Big)^{n}+b],$ (5)
@@
Figure 2: Percent error of $K(k)$ and $E(k)$ from various simple analytical forms
@@
Figure 3: Error of $K(k)$ and $E(k)$ from various mathematical inequalities
Fig. 3 (left) shows the performance of the upper and lower bounds for the complete elliptic integral $K(k)$. By resolving conjectures on inequalities between the $K(k)$ and simple logarithmic expressions, Qiu and coauthors derived sharp two-sided bounds for $K(k)$ [17]. These bounds are asymptotically exact as $k\to 0$ and as $k\to 1$. The Qiu et al. bounds for $k\in(0,1)$ are:
$\ln\bigg(\frac{4}{\sqrt{1-k^{2}}}+(e^{\pi/2}-4)\sqrt{1-k^{2}}\bigg)<K(k)<\ln\bigg(\frac{4}{\sqrt{1-k^{2}}}+e^{\pi/2}-4\bigg).$ (6)
Fig. 3 (right) illustrates the sharp bounds for the complete elliptic integral $E(k)$. Yang and coauthors derived sharp lower and upper bounds for $E(k)$ in terms of elementary functions [20, 22]. The upper bound is also exact in the two limiting cases. In the inset, the approximation $E(k)$ in this work intersects the upper bound at $k\approx 0.964$. This crossover suggests an interesting property that may merit further investigation within the framework of inequality analysis. The Yang et al. bounds for $k\in(0,1)$ are of the forms:
$\frac{7\pi}{22}\frac{\,\,1-(1-k^{2})^{11/8}}{1-(1-k^{2})^{7/8}}<E(k)<\frac{\,\,1-(1-k^{2})^{\pi/(2\pi-4)}}{1-(1-k^{2})^{1/(\pi-2)}}.$ (7)
@@ 3.1 Nonlinear Pendulum
Fig. 4 shows the phase space of an energy-conserving nonlinear pendulum with physical parameters $\sqrt{L/g}=1\,\text{second}$. The $\theta(t)$ are as follows:
$\omega_{m}<\omega_{c}:$ $\quad\theta(t)=\sum_{n\;\text{odd}}a_{n}\sin(n\Omega_{0}t+n\delta),$ (8)
$\omega_{m}=\omega_{c}:$ $\quad\theta(t)=2\arcsin\big[\tanh(\sqrt{\tfrac{g}{L}}t+\delta)\,\big],$ (9)
$\omega_{m}>\omega_{c}:$ $\quad\theta(t)=\Omega_{0}t+\delta+\sum_{n\geq 1}b_{n}\sin(n\Omega_{0}t+n\delta).$ (10)
@@
Specifically, the period $T$ for both classes of motion are ($k=\sqrt{L/g}\,\omega_{m}/2$):
Swinging: $\quad T=\sqrt{\frac{L}{g}}\,4\,K(k),$ (11)
Spinning: $\quad T=\sqrt{\frac{L}{g}}\,\frac{2}{k}\,K(1/k).$ (12)
@@ 4 Theoretical Method
Complete asymptotic behaviors of $K(k)$ and $E(k)$ are rarely available in mathematical handbooks. Fortunately, they are documented in the NIST Digital Library of Mathematical Functions (DLMF) [29] and serve as the foundation for the proposed formulae in this work. The complete asymptotic behaviors are given in Ref. [29, Eq.19.12.1-2]:
$K(k)$ $=\sum_{m=0}^{\infty}\frac{{(\frac{1}{2})}_{m}{(\frac{1}{2})}_{m}}{m!m!}{k^{\prime}}^{2m}\Big[\ln(\frac{1}{k^{\prime}})+d(m)\Big],$ (14)
$E(k)$ $=1+\frac{1}{2}\sum_{m=0}^{\infty}\frac{{(\frac{1}{2})}_{m}{(\frac{3}{2})}_{m}}{{(2)}_{m}m!}{k^{\prime}}^{2m+2}\Big[\ln(\frac{1}{k^{\prime}})+d(m)-\frac{1}{(2m+1)(2m+2)}\Big].$ (15)
where $k^{\prime}=\sqrt{1-k^{2}}$. The parenthesis with subscript, for example ${(\frac{1}{2})}_{m}$, is Pochhammer’s symbol. The auxiliary function $d(m)$ has a recursive relation $d(m+1)=d(m)-\frac{2}{(2m+1)(2m+2)}$, and $d(0)=\ln 4$. To find a balance between simplicity and accuracy, only the first term $m=0$ is kept. They become:
$\text{Asymptotic limit:}\qquad K(k)$ $\approx\ln\Big[\frac{4}{\sqrt{1-k^{2}}}\Big],$ (16)
$\text{Asymptotic limit:}\qquad E(k)$ $\approx 1+\frac{1-k^{2}}{2}\ln\Big[\frac{4/\!\sqrt{e}}{\sqrt{1-k^{2}}}\Big].$ (17)
@@
The following form is proposed:
$K(k)\approx\frac{1}{n}\ln[\Big(\frac{4}{\sqrt{1-k^{2}}}\Big)^{n}+b].$ (18)
@@
The following form is proposed:
$E(k)\approx 1+\frac{1-k^{2}}{2n}\ln[\Big(\frac{4/\!\sqrt{e}}{\sqrt{1-k^{2}}}\Big)^{n}+b].$ (19)